Net force on q1?

Please help me to complete the answer ASAP!

Thanks

Coulomb's Law: Example Net force on q1? q1 q2 q3 2
Coulomb's Law: Example Net force on q1? q1 q2 q3 2
Answer #1

Force on a charge q due to q' when separated by a distance r.

\vec{F} = \frac{1}{4\pi \epsilon_o}[\frac{qq'}{r^2}]\hat{r}

So Force on q1 due to q3 is given by:

\vec{F}_{13} = -\frac{1}{4\pi \epsilon_o}[\frac{(2)(2)}{(0.04)^2}]\hat{x} = -2.25\times 10^{13}\hat{x}N

And Force on q1 due to q2 is given by:

\vec{F}_{12} = -\frac{1}{4\pi \epsilon_o}[\frac{(2)(2)}{(0.04)^2}](cos60\degree \hat{x} + sin60\degree \hat{y}] = -1.125\times 10^{13}\hat{x}N - 1.95\times 10^{13}\hat{y}N

So the net force on q1 will be:

\vec{F}_{} = -(1.125 + 2.25)\times 10^{13}\hat{x}N - 1.95\times 10^{13}\hat{y}N

\vec{F}_{} = -3.375 \times 10^{13}\hat{x}N - 1.95\times 10^{13}\hat{y}N